I was absent due to the flu so i'm at a lost on what to do.  Ignore 1 continued.


L continued: Suppose you connect another capacito
Answer #1

Capacitor discharge (charge decay): Q = Qoe-(t/RC)

so at t = 0, Q = Q0 = cv = 125x10^-7C


b) i = V/R = 6.25mA


c)Q=Q0E^(-t/RC)


Q/Q0 = 0.01

t = -RCln(0.01)


t = 921x10^-5s